OpenStudy (dls):

The graph between linear momentum and time acting on body is given by the equation P=10+5t.If the mass of the body is 1Kg,find the acceleration of the body.

OpenStudy (dls):

\[\LARGE P=MV\] \[\LARGE P=V\] \[\LARGE 10+5t=V\] \[\LARGE dV=(5+10t)dt\] \[\LARGE \int\limits_{0}^{V} dV=\int\limits_{0}^{t}(5+10t)dt\] Is it how it is supposed to be done?just tried..

hartnn (hartnn):

how u got dV = 5+10 t dt ?

OpenStudy (shubhamsrg):

-_-

hartnn (hartnn):

also, acceleration is just dV/dt

OpenStudy (dls):

oh sorry i took it wrong

hartnn (hartnn):

V = 10+5t acceleration =dV/dt=... ?

OpenStudy (dls):

5

hartnn (hartnn):

that would be correct.

OpenStudy (dls):

lol :|